Cephalexin 25 mg/kg/day PO divided q 6h is ordered for a child with a beta-hemolytic streptococcal infection. The child weighs 20 kg. Your bottle of Cephalexin is labeled 250 mg / 5 mL. How many mL will you give the child every 6 hours?
Answer #1

So for calculation first we need to list down the points like as

Child's body weight=20 kg

Ordered dose of cephalexin 25 mg/kg/day p/o divided q6h ( means every 6 hrs in total 24 hrs that means 4 times a day)

According to child's body weight the dose will be

25 mg per 1 kg body weight

So required dose of mg will be of 20 kg body wt= 25×20=500 mg

Hence the the dose is 500 mg p/o in 24 hrs

Now we have to divide the dose into 4 times equally

That will be=500/4=125 mg

And i have bottle of cephalexin labeled250mg/5ml

5 ml contain 250 mg

1 ml contain 250/5mg

Answer 1ml =50 mg

Now we need to know how many ml contain 125 mg

50 mg in 1 ml

1 mg in 1/50

125 mg in 125/50

The answer is =2.5 ml

So 2.5 ml i will give the child every 6 hours